For the reaction





3 H2(g) + N2(g) ----------> 2 NH3(g)


what mass of NH3(g) is produced when 2.15 L of H2(g) measured at STP reacts?



1
Expert's answer
2022-05-19T00:58:03-0400

3 H2 + N2 => 2 NH3

1. under normal conditions, 1 mole of gaseous substance has a volume of 22.4 liters. Based on the above, we find the amount of substance (mol) in 2.15 liters of H2. for this we divide the given volume by 22.4:

n= 2.15 / 22.4 ≈ 0.096 mol

2. According to the above reaction, 3 moles of H2 react to form 2 moles of NH3. Using this, we find how much NH3 is formed when 0.096 mol H2 reacts:

3mol H2 —> 2mol NH3

0.096mol H2 —> x NH3

X= 0.096 * 2 / 3 = 0.064 mol NH3

3. Using the formula m = M * n, we find the mass of NH3:

m — mass of matter

M — molecular mass of matter (for NH3 =17gr/mol)

n — amount of substance

m = n * M = 0.064 * 17 = 1.088 gr

ANSWER: 1,088 gr NH3